OpenStudy (anonymous):

Solve by the substitution method 6x+5y=3 x=16-6y

OpenStudy (anonymous):

Minus the 6y on the 2nd equation

OpenStudy (anonymous):

You will get 6x+5y=3 x+6y=16 Then you will have to get rid of one of the variables

OpenStudy (anonymous):

Say you want to get rid of the x so will have to multiply -6 on the second equation then you will get -31y=-93

OpenStudy (anonymous):

Then the y=3 You plug the 3 into the 2nd equation x+6(3)=16 x+18=16 x=-2

OpenStudy (anonymous):

So x=-2 & y=3